# The Coordination Gap: Alternation Metrics for Temporal Dynamics in Multi-Agent Battle of the Exes

### 06 Mar 2026 v1

@@ 4 Proposed Assessment Framework
Traditional metrics fail to capture temporal structure—how access to high-payoff states is distributed across agents over time. We introduce Alternation Measures (ALT) that explicitly quantify who succeeds, when, and in what sequence.
@@ 4.1 Perfect Alternation as Reference
Perfect Alternation (PA) is a coordination regime where each of $n$ agents reaches the highest-payoff state exactly once within every block of $n$ episodes, in any order. Stricter order-sensitive variants (e.g., requiring each agent to occupy each rank equally often across batches) are possible but not imposed here. PA can be used descriptively, normatively, or prescriptively depending on the application; here it serves only as a reference regime against which observed coordination can be compared Papadopoulos (2020); Papadopoulos and Sanchez-Fibla (2021). Human participants in BoE experiments sometimes converge to PA-like patterns Hawkins and Goldstone (2016), motivating its use as a meaningful boundary condition. PA is Pareto-optimal under strict rotation, and in the two-agent ballistic case it constitutes a Nash equilibrium; detailed equilibrium links and the three motivating reasons (empirical observability, temporal sensitivity, metric calibration) are provided in Supplementary Section S4.
@@ 4.2 ALT Framework
ALT measures satisfy: (i) boundedness in $[0,1]$; (ii) $\mathrm{ALT}=1$ under PA; (iii) monotonic degradation under monopolization or ties; (iv) sensitivity to exclusivity; (v) permutation invariance Papadopoulos and Sanchez-Fibla (2021); Papadopoulos et al. (2025). For this study, order within a batch is not penalized; any permutation is treated as ideal. Full symbol definitions and computation details are provided in Supplementary Section S4.
@@ 4.3 Six ALT Variants
Emphasizes exclusivity—episodes with single winners:
@@
CALT serves as the primary metric due to its balanced treatment of exclusivity and ties. FALT provides a tolerant baseline, EALT isolates exclusivity, and AALT sets a strict diagnostic bound. ALT measures complement—rather than replace—traditional metrics by capturing temporal coordination invisible to outcome-based evaluation.
@@ 5.5.1 Random Baseline Performance
For 2-agent systems, random policies achieve Efficiency = 0.818 and Fairness = 0.972—values that would typically be interpreted as indicating successful coordination. Even with 5 agents, random policies maintain Efficiency $\geq 0.532$ and Fairness $\geq 0.927$. These high traditional metric values persist despite the complete absence of any learned coordination strategy.
@@
Figure 1 visualizes this critical comparison, showing that Q-learning agents (blue bars) consistently achieve lower CALT values than random baselines (orange bars) across all agent configurations—a finding that challenges assumptions about learned coordination in this class of multi-agent systems. We emphasize that this result does not generalize to all MARL algorithms, but highlights a failure mode of independent tabular learners under temporal coordination demands. We report CALT explicitly as the primary alternation metric; other ALT variants are summarized in Table 4.
@@ 5.5.3 Perfect Alternation Equivalent Analysis
Performance drops sharply from 56.8% (2 agents) to 25.0% (5 agents) and remains near $\sim$22% for 8–10 agents, demonstrating that coordination difficulty scales sharply with system size. Notably, 10 Q-learning agents achieve coordination equivalent to only 2.19 perfectly alternating agents—barely one-fifth of the agents in the system exhibiting coordinated behavior.
@@ 6.1 Why Does Q-Learning Fail at Coordination?
Our finding—that independent Q-learning agents perform below random baselines—demands explanation. We identify four interrelated factors:
@@
2. Non-Stationary Opponent Dynamics: From each agent’s perspective, other agents’ evolving policies violate assumptions underlying standard convergence results for Q-learning. As opponents explore different strategies during early training, the environment appears stochastic and non-stationary. Once all agents simultaneously cease exploration (synchronized epsilon decay at 75% of episodes), they may converge to a suboptimal coordination regime (Nash-like, locally stable) where no single agent can unilaterally improve by alternating—a classic coordination failure Masiliūnas (2019). This synchronized decay may also contribute to mid-range bumps in EALT/qEALT: partial exploration increases exclusive wins before late-stage exploitation drives more simultaneous arrivals and ties.
@@
4. Tragedy of the Learning Commons: Each agent learns to exploit opportunities for $r_{high}$ greedily, maximizing individual expected return. Without global coordination mechanisms, this individual rationality leads to collective irrationality—agents interfere with each other’s learning trajectories, producing worse outcomes than random policies that at least distribute wins stochastically.
These factors compound as agent count increases, explaining the sharp drop in PA-equivalent performance from 56.8% (2 agents) to 25.0% (5 agents) and the low plateau around $\sim$22% for 8–10 agents. The coordination problem is not merely difficult—it is poorly aligned with the assumptions of independent tabular reinforcement learning.
@@ 6.2 Coordination Breakdown as Strategic Regime Shift
The performance drop from 56.8% (2 agents) to $\sim$22% (8–10 agents) reflects a fundamental shift in the equilibrium structure of multi-agent games: as agent count increases, the strategic complexity of achieving coordinated alternation grows sharply, and independent learners fail to discover equilibria supporting fair turn-taking Pikovsky (2009); Maruta and Okada (2012); Masiliūnas (2019); Zhu et al. (2020); Shan et al. (2019); He et al. (2015); Tian et al. (2026). Below a certain agent threshold, residual coordination persists through partial turn-taking patterns; beyond this threshold, individual rationality leads to collective irrationality—agents interfere with each other’s strategies, producing coordination outcomes qualitatively close to random dynamics.
The finding that 10 Q-learning agents coordinate as effectively as only 2.19 perfectly alternating agents constitutes a notable finding, not merely a negative result. It identifies a regime in which coordination breaks down in decentralized multi-agent systems—information that remains entirely hidden when using traditional metrics whose ranges can appear deceptively high (e.g., Reward Fairness up to 0.993 and Efficiency up to 0.677). Such breakdowns, where individual optimization undermines collective welfare, are central concerns in game theory, mechanism design, and the study of social dilemmas.
@@ 6.3 Implications for Random Baseline Methodology
In prior work on the Battle of the Exes and related repeated coordination games, random-policy baselines are not consistently reported as a standard reference, and in several cases the presence or absence of such baselines is not discussed explicitly Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023). This can make it difficult to assess whether observed coordination reliably exceeds chance-level behavior, particularly when outcome-based metrics take deceptively high values even under random play.
@@ 7 Conclusions
This study introduces Perfect Alternation (PA) as a reference coordination regime and limiting benchmark for temporal turn-taking, alongside six novel ALT metrics and a regression-based benchmarking framework enabling interpretable quantification of coordination quality. Through comprehensive Q-learning experiments with rigorous random policy baselines—a methodological practice not consistently reported in prior BoE-related studies—we reveal fundamental limitations in both traditional evaluation metrics and learned coordination strategies Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023).
@@ 7.1 Key Findings
1. Traditional Metrics Severely Fail to Detect Coordination Failure. Across all configurations, Q-learning agents exhibit Reward Fairness in the range 0.49–0.993, Efficiency 0.054–0.677, and TT Fairness 0.699–0.998—values that can appear deceptively high in aggregate. However, ALT metrics reveal the opposite: Relative Change reaches as low as -81% (qEALT), with only one small positive at +5.4% for 10-agent qEALT, demonstrating that learned policies perform worse than random baselines under ALT evaluation. This dichotomy exposes a fundamental inadequacy in outcome-based fairness measures, which cannot distinguish monopolistic, random, or alternating resource access patterns in temporal multi-agent scenarios.
2. Q-Learning Produces Systematic Under-Coordination in Multi-Agent BoE. With only one negligible exception (10-agent qEALT at +5.4%), all coordination scores are negative. This systematic underperformance relative to random baselines (to our knowledge, not previously established in BoE literature) indicates that independent Q-learning agents do not merely fail to achieve coordination—they learn policies that do not promote alternation. The coordination deficit is substantial across sizes (CALT coordination scores: -33.7% for 2 agents, -74.3% for 5 agents, -56.6% for 10 agents), with AALT and qEALT confirming this pattern across the full ALT strictness spectrum, suggesting that tabular Q-learning’s inability to model joint action spaces becomes increasingly problematic as system complexity grows.
3. Coordination Difficulty scales sharply with Agent Count. PA-equivalent analysis shows a steep drop from 56.8% of perfect coordination (2 agents) to 25.0% (5 agents) and a low plateau around $\sim$22% for 8–10 agents. This indicates that coordination challenges compound non-linearly as competitive pressure increases. The finding that 10 agents achieve coordination equivalent to only 2.19 perfectly alternating agents—representing about one-fifth of the system—quantifies the severity of the multi-agent coordination problem in decentralized learning scenarios.
@@ 7.2 Theoretical and Methodological Contributions
ALT Metrics address critical gaps in existing fairness literature by integrating distributional fairness (min/max ratios) with temporal fairness (rotation periodicity). The six metrics offer tunable strictness—from FALT’s lenient fractional evaluation to AALT’s exclusive-win requirement—enabling researchers to select appropriate sensitivity levels for different experimental phases or application domains. Critically, these metrics properly expose coordination failure that traditional measures mask, as demonstrated by our random baseline comparisons.
@@
While random baselines are widely used in other areas of empirical evaluation, they have not been consistently adopted as standard practice in studies of the Battle of the Exes and related coordination games, limiting comparability across works and complicating claims of emergent coordination Hawkins and Goldstone (2016); Puig Camps (2018); Gasparrini and Sanchez-Fibla (2018); Freire et al. (2020, 2023).
@@ 7.3 Implications, Limitations, and Future Directions
By establishing rigorous baselines, developing interpretable benchmarking methodologies, and exposing limitations in conventional evaluation practices, this work advances both theoretical understanding and practical assessment of multi-agent coordination. More broadly, our findings contribute to the social choice and welfare measurement literature by demonstrating that distributional fairness criteria—widely used in fair division and cooperative game theory—can be fundamentally misleading in repeated interaction settings where the temporal sequence of allocations matters as much as their aggregate distribution. The result that Q-learning agents perform below random baselines—despite appearing successful by traditional measures—underscores the necessity of domain-appropriate metrics and the dangers of relying on outcome fairness alone when procedural fairness matters.

### 10 Mar 2026 v2

@@ 4 Proposed Assessment Framework
Traditional metrics fail to capture temporal structure (how access to high-payoff states is distributed across agents over time). We introduce Alternation Measures (ALT) that explicitly quantify who succeeds, when, and in what sequence.
@@ 4.1 Perfect Alternation as Reference
Perfect Alternation (PA) is a coordination regime where each of $n$ agents reaches the highest-payoff state exactly once within every block of $n$ episodes, in any order. Stricter order-sensitive variants (e.g., requiring each agent to occupy each rank equally often across batches) are possible but not imposed here. PA can be used descriptively, normatively, or prescriptively depending on the application; here it serves only as a reference regime against which observed coordination can be compared (Papadopoulos, 2020; Papadopoulos and Sanchez-Fibla, 2021). Human participants in BoE experiments sometimes converge to PA-like patterns (Hawkins and Goldstone, 2016), motivating its use as a meaningful boundary condition. PA is Pareto-optimal under strict rotation, and in the two-agent ballistic case it constitutes a Nash equilibrium; detailed equilibrium links and the three motivating reasons (empirical observability, temporal sensitivity, metric calibration) are provided in Supplementary Section S4.
@@ 4.2 ALT Framework
ALT measures satisfy: (i) boundedness in $[0,1]$; (ii) $\mathrm{ALT}=1$ under PA; (iii) monotonic degradation under monopolization or ties; (iv) sensitivity to exclusivity; (v) permutation invariance (Papadopoulos and Sanchez-Fibla, 2021; Papadopoulos et al., 2025). For this study, order within a batch is not penalized; any permutation is treated as ideal. Full symbol definitions and computation details are provided in Supplementary Section S4.
@@ 4.3 Six ALT Variants
Emphasizes exclusivity (episodes with single winners):
@@
CALT serves as the primary metric due to its balanced treatment of exclusivity and ties. FALT provides a tolerant baseline, EALT isolates exclusivity, and AALT sets a strict diagnostic bound. ALT measures complement (rather than replace) traditional metrics by capturing temporal coordination invisible to outcome-based evaluation.
@@ 5.5.1 Random Baseline Performance
For 2-agent systems, random policies achieve Efficiency = 0.818 and Fairness = 0.972, values that would typically be interpreted as indicating successful coordination. Even with 5 agents, random policies maintain Efficiency $\geq 0.532$ and Fairness $\geq 0.927$. These high traditional metric values persist despite the complete absence of any learned coordination strategy.
@@
Figure 1 visualizes this critical comparison, showing that Q-learning agents (blue bars) consistently achieve lower CALT values than random baselines (orange bars) across all agent configurations, a finding that challenges assumptions about learned coordination in this class of multi-agent systems. We emphasize that this result does not generalize to all MARL algorithms, but highlights a failure mode of independent tabular learners under temporal coordination demands. We report CALT explicitly as the primary alternation metric; other ALT variants are summarized in Table 4.
@@ 5.5.3 Perfect Alternation Equivalent Analysis
Performance drops sharply from 56.8% (2 agents) to 25.0% (5 agents) and remains near $\sim$22% for 8–10 agents, demonstrating that coordination difficulty scales sharply with system size. Notably, 10 Q-learning agents achieve coordination equivalent to only 2.19 perfectly alternating agents, barely one-fifth of the agents in the system exhibiting coordinated behavior.
@@ 6.1 Why Does Q-Learning Fail at Coordination?
Our finding that independent Q-learning agents perform below random baselines demands explanation. We identify four interrelated factors:
@@
2. Non-Stationary Opponent Dynamics: From each agent’s perspective, other agents’ evolving policies violate assumptions underlying standard convergence results for Q-learning. As opponents explore different strategies during early training, the environment appears stochastic and non-stationary. Once all agents simultaneously cease exploration (synchronized epsilon decay at 75% of episodes), they may converge to a suboptimal coordination regime (Nash-like, locally stable) where no single agent can unilaterally improve by alternating, a classic coordination failure (Masiliūnas, 2019). This synchronized decay may also contribute to mid-range bumps in EALT/qEALT: partial exploration increases exclusive wins before late-stage exploitation drives more simultaneous arrivals and ties.
@@
4. Tragedy of the Learning Commons: Each agent learns to exploit opportunities for $r_{high}$ greedily, maximizing individual expected return. Without global coordination mechanisms, this individual rationality leads to collective irrationality, as agents interfere with each other’s learning trajectories, producing worse outcomes than random policies that at least distribute wins stochastically.
These factors compound as agent count increases, explaining the sharp drop in PA-equivalent performance from 56.8% (2 agents) to 25.0% (5 agents) and the low plateau around $\sim$22% for 8–10 agents. The coordination problem is not merely difficult; it is poorly aligned with the assumptions of independent tabular reinforcement learning.
@@ 6.2 Coordination Breakdown as Strategic Regime Shift
The performance drop from 56.8% (2 agents) to $\sim$22% (8–10 agents) reflects a fundamental shift in the equilibrium structure of multi-agent games: as agent count increases, the strategic complexity of achieving coordinated alternation grows sharply, and independent learners fail to discover equilibria supporting fair turn-taking (Pikovsky, 2009; Maruta and Okada, 2012; Masiliūnas, 2019; Zhu et al., 2020; Shan et al., 2019; He et al., 2015; Tian et al., 2026). Below a certain agent threshold, residual coordination persists through partial turn-taking patterns; beyond this threshold, individual rationality leads to collective irrationality, as agents interfere with each other’s strategies, producing coordination outcomes qualitatively close to random dynamics.
The finding that 10 Q-learning agents coordinate as effectively as only 2.19 perfectly alternating agents constitutes a notable finding, not merely a negative result. It identifies a regime in which coordination breaks down in decentralized multi-agent systems, information that remains entirely hidden when using traditional metrics whose ranges can appear deceptively high (e.g., Reward Fairness up to 0.993 and Efficiency up to 0.677). Such breakdowns, where individual optimization undermines collective welfare, are central concerns in game theory, mechanism design, and the study of social dilemmas.
@@ 6.3 Implications for Random Baseline Methodology
In prior work on the Battle of the Exes and related repeated coordination games, random-policy baselines are not consistently reported as a standard reference, and in several cases the presence or absence of such baselines is not discussed explicitly (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018; Freire et al., 2020, 2023). This can make it difficult to assess whether observed coordination reliably exceeds chance-level behavior, particularly when outcome-based metrics take deceptively high values even under random play.
@@ 7 Conclusions
This study introduces Perfect Alternation (PA) as a reference coordination regime and limiting benchmark for temporal turn-taking, alongside six novel ALT metrics and a regression-based benchmarking framework enabling interpretable quantification of coordination quality. Through comprehensive Q-learning experiments with rigorous random policy baselines (a methodological practice not consistently reported in prior BoE-related studies), we reveal fundamental limitations in both traditional evaluation metrics and learned coordination strategies (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018; Freire et al., 2020, 2023).
@@ 7.1 Key Findings
1. Traditional Metrics Severely Fail to Detect Coordination Failure. Across all configurations, Q-learning agents exhibit Reward Fairness in the range 0.49–0.993, Efficiency 0.054–0.677, and TT Fairness 0.699–0.998, values that can appear deceptively high in aggregate. However, ALT metrics reveal the opposite: Relative Change reaches as low as -81% (qEALT), with only one small positive at +5.4% for 10-agent qEALT, demonstrating that learned policies perform worse than random baselines under ALT evaluation. This dichotomy exposes a fundamental inadequacy in outcome-based fairness measures, which cannot distinguish monopolistic, random, or alternating resource access patterns in temporal multi-agent scenarios.
2. Q-Learning Produces Systematic Under-Coordination in Multi-Agent BoE. With only one negligible exception (10-agent qEALT at +5.4%), all coordination scores are negative. This systematic underperformance relative to random baselines (to our knowledge, not previously established in BoE literature) indicates that independent Q-learning agents do not merely fail to achieve coordination; they learn policies that do not promote alternation. The coordination deficit is substantial across sizes (CALT coordination scores: -33.7% for 2 agents, -74.3% for 5 agents, -56.6% for 10 agents), with AALT and qEALT confirming this pattern across the full ALT strictness spectrum, suggesting that tabular Q-learning’s inability to model joint action spaces becomes increasingly problematic as system complexity grows.
3. Coordination Difficulty scales sharply with Agent Count. PA-equivalent analysis shows a steep drop from 56.8% of perfect coordination (2 agents) to 25.0% (5 agents) and a low plateau around $\sim$22% for 8–10 agents. This indicates that coordination challenges compound non-linearly as competitive pressure increases. The finding that 10 agents achieve coordination equivalent to only 2.19 perfectly alternating agents (representing about one-fifth of the system) quantifies the severity of the multi-agent coordination problem in decentralized learning scenarios.
@@ 7.2 Theoretical and Methodological Contributions
ALT Metrics address critical gaps in existing fairness literature by integrating distributional fairness (min/max ratios) with temporal fairness (rotation periodicity). The six metrics offer tunable strictness (from FALT’s lenient fractional evaluation to AALT’s exclusive-win requirement), enabling researchers to select appropriate sensitivity levels for different experimental phases or application domains. Critically, these metrics properly expose coordination failure that traditional measures mask, as demonstrated by our random baseline comparisons.
@@
While random baselines are widely used in other areas of empirical evaluation, they have not been consistently adopted as standard practice in studies of the Battle of the Exes and related coordination games, limiting comparability across works and complicating claims of emergent coordination (Hawkins and Goldstone, 2016; Puig Camps, 2018; Gasparrini and Sanchez-Fibla, 2018; Freire et al., 2020, 2023).
@@ 7.3 Implications, Limitations, and Future Directions
By establishing rigorous baselines, developing interpretable benchmarking methodologies, and exposing limitations in conventional evaluation practices, this work advances both theoretical understanding and practical assessment of multi-agent coordination. More broadly, our findings contribute to the social choice and welfare measurement literature by demonstrating that distributional fairness criteria—widely used in fair division and cooperative game theory—can be fundamentally misleading in repeated interaction settings where the temporal sequence of allocations matters as much as their aggregate distribution. The result that Q-learning agents perform below random baselines (despite appearing successful by traditional measures) underscores the necessity of domain-appropriate metrics and the dangers of relying on outcome fairness alone when procedural fairness matters.